Presented below are two independent situations:

(a) On January 1, 2020, Pina Inc. purchased land that had an assessed value of $373,000 at the time of purchase. A $530,000, zero-interest-bearing note due January 1, 2023, was given in exchange. There was no established exchange price for the land, nor a ready fair value for the note. The interest rate charged on a note of this type is 12%.

Determine at what amount the land should be recorded at January 1, 2020, and the interest expense to be reported in 2020 related to this transaction. (Round answers to 0 decimal places, e.g. 38,548.)

Land to be recorded at January 1, 2020 $
Interest expense to be reported $


(b) On January 1, 2020, Grouper Furniture borrowed $3,600,000 (face value) from Sinise Co., a major customer, through a zero-interest-bearing note due in 4 years. Because the note was zero-interest-bearing, Grouper Furniture agreed to sell furniture to this customer at lower than market price. A 10% rate of interest is normally charged on this type of loan.

Prepare the journal entry to record this transaction. (Round answers to 0 decimal places, e.g. 38,548. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Account Titles and Explanation

Debit

Credit


Determine the amount of interest expense to report for 2020. (Round answer to 0 decimal places, e.g. 38,548.)

Interest expense to be reported for 2020 $
Answer #1

Answer :-

a)On January 1, 2020, Pina Inc. purchased land that had an assessed value of $373,000.

Zero-interest-bearing note on January 1,2023 - $530,000 was given in exchange.

Now we calculate the Present value of Zero-interest-bearing.

PV factor for 3 year at 12% = 1/(1 +r )^n

PV factor for 3 year at 12%= 1/(1 + 12% ) ^3

PV factor for 3 year at 12%= 0.7118

Present value of Zero-interest-bearing note = Future Value of Zero interest bearing notes × PV factor

Present value = $530,000 × 0.7118

Present value = $377,254

Land to be recorded at January 1, 2020 = Present value of Zero-interest-bearing note

Land to be recorded at January 1, 2020 = $377,254

The interest expense = Land value on January 1,2020 × Interest rate

Interest rate = 12%

The interest expense = $377,254 × 12%

Interest expense = $ 45,270.48

b)On January 1, 2020, Grouper Furniture borrowed $3,600,000 (face value) from Sinise Co.

Journal entries :-

Particular Debit Credit
Cash A/c . Dr . $3,600,000
Discount on Note payableA/c Dr. $1,141,200
To Notes Payable $3,600,000
To Unearned revenue (Note -1) $1,141,200

Note 1: -

Future value zero-interest-bearing note due in 4 years - $3,600,000

Present value of Zero-interest-bearing note = Future Value of Zero interest bearing notes × PV factor

PV factor for 4 year at 10% = 1/(1 +r )^n

PV factor for 4 year at 10%= 1/(1 + 10% )⁴

PV factor for 4 year at 10%= 0.6830

Present value of note - $3,600,000 × 0.6830

Present value of note - $2,458,800

Unearned revenue = Future value - Present value of note

Unearned revenue = $3,600,000 - $2,458,800

Unearned revenue = $1,141,200

.

Interest Expense = Present value of note × Interest rate

Interest rate = 10%

Interest expense= $2,458,800 × 10%

Interest expense= $245,880